Fermi National Laboratory


Accelerator Update

Wednesday April 21
The day shift began with Operations monitoring store 3423 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with MiniBooNE taking beam, and with the Meson MCenter experiment on access.

At 11:05 AM, MCenter came out of access.

Problems with a Pbar device (PMAG) halted stacking at 11:33 AM. An expert investigated and initially discovered a shorted SCR, but later they discovered a problem with the magnet and it had to be replaced.

The Tevatron (TeV) suffered a sector E1 quench at 11:52 AM due to a CAMAC crate power supply failure.

Since Pbar was off, the Booster access was moved up to 12:32 AM. Pbar experts also accessed their ring at this time to look at the pulsed magnet.

Safety system technicians accessed the Linac at 12:54 AM.

Booster experts came out of access at 2:08 PM.

Operations established beam to MiniBooNE and to the Main Injector (MI) dump at 2:51 PM.

TeV studies began at 2:56 PM and Booster studies began a few minutes later.

Operations put the TeV through a dry squeeze at 5:36 PM.

At 7:02 PM, Operators discovered a problem with equipment related to Debuncher RF station (DRF1-5). After trying various fixes and consulting with an expert they discovered and replaced a blown fuse.

At 10:09 PM, a Pbar expert began a thirty-minute Accumulator RF study.

Operations established beam to MCenter at 10:35 PM.

Thursday April 22
The midnight shift began with technicians replacing the Pbar pulsed magnet, with Operations monitoring the Recycler stash, with experts conducting studies in the TeV and MI, and with MiniBooNE and MCenter taking beam.

Techs in Pbar reported at 5:15 AM that they were having problems with running the pulsed magnet. Operators called in another expert.

A Booster power supply (GMPS) tripped off at 5:41 AM due to high vacuum pressure. The Booster chopper located in the Linac 400 MeV area needed to be power cycled to resume beam at 6:07 AM.

At 7:53 AM, techs reported that the test pulsing of the Pbar pulsed magnet had finished.

The MCenter experiment stopped taking beam at 8:45 AM to work on equipment.

Pbar began stacking at 9:53 AM.

TeV studies ended at 1:56 PM. Operators had the Collider Detector at Fermilab (CDF) and D-zero (D0) ready for access by 2:12 PM. Both experiments were out of access by 3:43 PM.

Operations put the TeV through a dry squeeze at 4:05 PM.

At 5 PM, Operations loaded the TeV with six bunches of protons to check the Tevatron Electron Lens (TEL) frequency. Operations aborted this beam at 5:24 PM.

Operations put the TeV through a wet squeeze at 5:41 PM.

Operations put the TeV into shot setup at 6:25 PM.

At 8:43 PM, Operations established store 3434 with an initial luminosity of 44.38 E30. Pbar resumed stacking.

Operators interlocked MCenter enclosures MC7 and MC8 at 10:47 PM.

Operations began preparing Recycler for stashing antiprotons.

Friday April 23
The midnight shift began with Operations monitoring store 3434 and a Recycler stash, with operators preparing Recycler and Pbar to stash more antiprotons, and with MiniBooNE taking beam.

At 12:44 AM, Pbar had completed its first shot to the Recycler, but ten minutes later a MI quadrupole power supply (M:Q109) tripped off.

Pbar resumed stacking at 12:56 AM.

At 1:14 AM, Operations established beam to MCenter.

CDF requested Operations to fix a beam problem at 5:28 AM. Operators discovered two collimators not in their proper position. Operators moved these collimators to their correct position at 7:19 AM, but it didn't fix the problem. CDF and D0 stopped taking data. Operators contacted an expert.

At 7:44 AM, Operations moved another collimator, suggested by the expert, into the beam and this solved the problem. CDF and D0 turned back on.

Plans for Today and this Weekend
The plans for today and this weekend are to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/23/2004   email Fermilab